This table shows 10 observations of a pair of variables (x,y). The variables x and y are positively correlated with a correlation coefficient of 0.977 Find the slope b, of the least squares regression line, y= ax + b, for these date.
____________________________________

X 1 2 3 4 5 6 7 8 9 11

Y 2 4 5 8 9 12 13 14 15 16
____________________________________

I TRY THIS: (16-2)= M(11-1),10M = 14, M = 1.4

BUT IT'S WRONG

Answer by rfer(16322)

m=(2-4)/(1-2)
m=-2/-1
m=2
---------------------
y=mx+b
2=2*2+b
2=4+b
b=-2
--------------------
y=2x-2